Here is a set of sample data

3 9 15 23 29
33 38 41 48 54
60 64 75 81 83
84 85 91 97


Identify the 5 number summary (min, Q1, median, Q3, max)
, , , ,
Answer #1

First , The data should be  arranged in ascending order

Minimum = 3

Q1 is the first quartile , under which 25% data lies

n= 19

Q1 is the (19+1)/ 4= 5th observation

Q1= 29

Median divides the data in two equal parts

Median is the (19+1)/2 = 10th observation

Median = 54

Q3 is the third  quartile , under which 75% data lies

Q1 is the 3(19+1)/ 4= 15th observation

Q3 = 83

Maximum = 97

Thus ,answer is

minimum = 3 , Q1 = 29 , median = 54 , Q3 = 83 . maximum = 97

3
9
15
23
29
33
38
41
48
54
60
64
75
81
83
84
85
91
97